OpenStudy (anonymous):

How can 1/5x-2=1/3x+8 be set up as a system of equations

OpenStudy (tkhunny):

Separate the left hand side from the right hand side and set both " = y".

OpenStudy (anonymous):

dang it i was about to say that

OpenStudy (anonymous):

okay i just did that

OpenStudy (anonymous):

what would the next step be

OpenStudy (anonymous):

@tkhunny go ahead bc i dont know the next step

OpenStudy (tkhunny):

There isn't a next step. Show your work and you are done.

OpenStudy (anonymous):

oh thats why i didnt know it lol

OpenStudy (anonymous):

5y − 5x = −10 3y − 3x = 24 5y − 5x = −10 3y + 3x = 24 5y + x = −10 3y + x = 24 5y − x = −10 3y − x = 24

OpenStudy (anonymous):

those are my choices im confused

OpenStudy (bee_see):